OpenStudy (anonymous):

The density of ice is 917 kg/m3, and the density of sea water is 1025 kg/m3. A swimming polar bear climbs onto a piece of floating ice that has a volume of 5.45 m3. What is the weight of the heaviest bear that the ice can support without sinking completely beneath the water?

OpenStudy (anonymous):

 \[v \times \rho(water)\times g =m(ice)\times g +M(maximum-mass)\times g\] \[v \times \rho(water)\times g =v \times \rho(ice) \times g +M(maximum- mass)\times g\] v is the volume of ice and M is the maximum mass of the polar bear..find M..